OpenStudy (anonymous):

The diameter of a circular garden is 100 metres.There is a 3m wide white flowers area running around its inside.Find its area and the cost of planting white flowers if cost of plating them is Rs.20 per square metre

OpenStudy (anonymous):

total area=62800 divide it by 3 and multiply with 20 you get the answer

OpenStudy (anonymous):

4186

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

radius of garden (r1) = 50m radius of garden including flowers area (r2) = 103/2 m Area of flower area = Area of garden including flowers area - Area of garden = \[\pi (r2)^{2}-\pi (r1)^2\]taking pi common, \[\pi [(r2)^{2}- (r1)^2]\] \[\pi (r2+r1)(r2-r1)\] =22/7 * 203/2 * 3/2 = 478.5 sq. m Cost of painting = Rs.20/sq. m Cost of painting 478.5sq. m = Rs20*478.5 = Rs 9570

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

oh!! well my answer was wrong i ddnt assumed a flower as an individual circle please consider the other1

OpenStudy (anonymous):

its allright!